OpenStudy (anonymous):

write an equation for the circle satisfying the given equations. radius 5, center (5,-3)

OpenStudy (experimentx):

(x-5)^2+(y-(-3))^2 = 5^2

OpenStudy (anonymous):

I got (x-5)^2 + (y+3)^2 =225.. I meant to say the radius was 15. sorry